# Improvement and generalization of ABCD method with Bayesian inference

### v1

@@ 1 Introduction
A more involved, but still brilliantly simple, data-driven modelling technique is the ABCD method [1, 2]. This approach consists in finding two independent observables that classify the background into four regions, such that the signal lies only in one of them. The method allows to predict the signal events in its corresponding region as long as its underlying hypotheses are satisfied. The method has been successfully implemented by the community and it can be found in many of the ATLAS [3] and CMS [4] public results, with some particular examples listed in Refs. [5, 6, 7, 8, 9]. In many applications, the method is adapted or extended to the particular features of the observable and backgrounds. The ABCD approach is quite useful because it reduces the impact of potential biases and inaccuracies in simulations, and because of its simplicity it only relies in finding the observables and regions that fulfill the required assumptions.
@@
In this work we propose to study an existing method from the area of Statistics, more specifically from Bayesian Machine Learning, that consists in disentangling many classes which are mixed in a given dataset. This procedure, known as a mixture model [13], considers that the dataset ${\bf X}$ is a realization of a probability density function with given parameters $\theta$. Using the Bayes Theorem, the method yields –most of the times numerically– a probability distribution for the parameters of the model, $p(\mbox{$\theta$}|{\bf X})$, given a collection of data ${\bf X}$, namely events at LHC, each containing the value of several observables measured in the same event. This distribution is called the posterior of the model. In contrast to ABCD, observables in the mixture model can have continuous distributions that, by definition, contain more information than discrete outputs. A mixture model as presented here assumes that each data point (event) belongs to only one of the possible classes.
The mixture model has among its parameters the expected fraction of each class in the dataset, and in particular for the class that corresponds to the signal, this fraction yields the expected amount of signal in the dataset, which is generally the sought-after unknown. In addition, the Bayesian inference also finds the posterior distribution for all the other parameters included in $\theta$ as well. This posterior refines the understanding of the physical system, since most of these parameters have a connection to the physics involved in the problem. Moreover, as it can be seen from the mathematical formulation of the mixture model, the number of classes is not fixed (while in the ABCD method there can only be signal and background), and the number of observables of each data point is also not limited to any number (while the ABCD method is restricted to two independent observables). For the purposes of simplicity and performance we also demand independence for the observables measured in each data point in the mixture model.  [^1] A more complex scenario with dependent observables could also be tackled through a mixture model as long as one has some prior-knowledge or clues about this dependence. We do not address this possibility in this work. Another recognizable feature of the mixture models is that they do not need control regions, and that there are no hard cuts to define regions nor hard-assignments to define whether a given event belongs to a specific class. Instead, each event is assigned a probability of belonging to each class.
@@
To compare both methods we use a toy problem whose basic properties are inspired by the di-Higgs production at LHC, with both Higgs states decaying to $b\bar{b}$, namely: $pp\to hh\to b\bar{b}b\bar{b}$. This process, being one of the best candidates to probe the nature of the Higgs boson, and one of the most relevant measurements to be expected by the LHC in the forthcoming years, has been studied with the ABCD method with some modifications in [8, 9]. It has several interesting features, such as a small cross-section, large backgrounds and, given the class, the presence of several independent or approximately independent observables, such as the four $b$-tagging scores [^2] There are systematic sources that may yield a small dependence between these scores. Although we ignore this dependence in the toy problem studied along is work, it should be taken into account in a more realistic study. of the jets and the two invariant masses of the corresponding paired jets. As a first step in the analysis of this process with Bayesian inference, we consider in this article a toy model that keeps some of its properties, as well as just two of its backgrounds, and analyse it with the ABCD method and with Bayesian inference. We study and quantify both performances and provide details on why the Bayesian analysis has very good perspectives.
@@ 2.1 The ABCD method: an overview
Where $N_{X}$ refers to the number of events in region $X$. $N_{A}(\mbox{background})$ refers to the number of background events in region $A$. This result is easily understood because if the observables have background distributions which are independent of each other, then one expects the ratio $N_{C}/N_{D}$ to be equal to $N_{A}({\mbox{background}})/N_{B}$. Observe that we assume that only region A contains signal. In the literature one can sometimes find that B is named as control region, and $N_{C}/N_{D}$ as a transfer function.
@@
As it can be appraised, the ABCD method is very clever and simple, and it has worked with excellent results and achievements in HEP, as discussed in the Introduction. However, if the hypotheses are not exactly satisfied, the predictions would deviate from their corresponding true values, as expected. Not only that, but also if the total number of events in B, C or D is small, then its Poisson fluctuation would propagate to the signal and background events predicted in A.
Therefore, it is also interesting to analyze some limitations in applying the ABCD method. As a first observation one should notice that the method uses hard cuts to define the regions; whereas mixture models can deal with all the information contained in the dataset. Using this information, applications in statistics [13] have shown improvements when using soft-assignments [14]. That is, each event has a given probability of being signal and a given probability of being background (in mixture models this is usually referred as responsibility [13]). As a second point, one should notice that the ABCD method is limited to two independent observables. This yields that when there are more independent observables then some of them are combined to reduce the number of observables, and/or the method is binned in the other observables. In these cases one computes many ABCD methods according to the binned output of the other observable(s). Also, if it is the case that there are many different backgrounds, then the ABCD method does not have the tools to exploit this knowledge beyond its formula in Eq. 1.
@@ 2.2 Improving and generalizing ABCD through Bayes inference techniques
We model this dataset from the mathematical point of view as the outcome of a probabilistic generative model depicted in the Graphical Model [13] in Fig. 2. We assume that, given the $n^{th}$ event, its class is determined through the sampling of a categorical latent random variable $\bf{z}_{n}$, which can take $K$ values and is not observed. Then, the $D$ observables in this event are sampled using the distributions from the class corresponding to the value of $\bf{z}_{n}$, and we assume that each observable has no dependence on the other $D-1$ observables once the class has been sampled. Henceforth, although $\bf{z}_{n}$ is not observed, the conditionally independent outcome of the $D$ observables provides information about the probability of the event to belong to each one of the classes. Moreover, the $D$ observables also provide information on their own distributions within each class. Bayesian inference provides a framework to exploit all this data.
In standard inference techniques [13], to write down the probability of a data point ${\bf x}_{n}$ and its corresponding latent variable $\bf{z}_{n}$, it is customary to use a 1-of-K representation for the latter. This implies that the latent variable for a given event is a vector with only one component equal to 1 and all other components 0. With this notation the probability for the $n^{th}$ data point becomes
$p({\bf x}_{n},{\bf z}_{n}|\theta)=p({\bf z}_{n}|{\bf\pi})\prod_{k=1}^{K}\,\left(\prod_{d=1}^{D}p(x_{nd}|\theta_{kd})\,\right)^{z_{n_{k}}},$ (2)
where $z_{n_{k}}$ is the $k^{th}$ component of ${\bf z}_{n}$ and $\theta$ represents the parameters of the model which can be split into the class fraction parameters ${\bf\pi}=\pi_{k=1,\dots K}$ and the parameters of the distributions for each of the $D$ random variables for the $k$-th class, $\theta_{kd}$. To shorten the notation, we will refer to the class-dependent distributions simply as $p({\bf x}_{n}|k)$. Having represented ${\bf z}_{n}$ as a 1-of-K vector, we observe how the product selects only the class with a non-zero value of $z_{n_{k}}$. Additionally, $\bf{z}_{n}$ follows a Categorical [^3] A categorical distribution is the usual multinomial distribution specialized to the case where there is only one draw. distribution with the same form
@@
$p({\bf x}_{n}|\theta)=\sum_{{\bf z}_{n}=\delta_{kk^{\prime}}}p({\bf x}_{n},{\bf z}_{n}|\theta),$ (4)
where with $\delta_{kk^{\prime}}$ we mean a vector with zeroes except in the $k$ index. Because each specific value of ${\bf z}_{n}$ selects one class $k$, the marginalized probability distribution takes the usual mixture model form
$p({\bf x}_{n}|\theta)=\sum_{k=1}^{K}\,\pi_{k}\,p({\bf x}_{n}|k).$ (5)
@@
$p({\bf X}|\theta)=\prod_{n=1}^{N}\,p({\bf x}_{n}|\theta)$ (6)
which is the basic ingredient to run Bayesian inference. We need to add a prior probability for the parameters of the model, $p(\theta)$, and one can in principle obtain the posterior $p(\theta|{\bf X})$. This usually can be achieved through a variety of numerical techniques, we describe in Section 3 the tools that we have used along this work.
@@
2 independent observables.
$D$ independent observables, with $D$ arbitrary.
2 components, signal and background.
$K$ components can make up the mixture model, the signal and $K-1$ backgrounds. $K$ arbitrary
Prior knowledge on signal and background distributions allows to define the four hard cut regions A, B, C and D, and to assume that the signal is only found in A. Then the method estimates how many signal events are expected in A.
Instead of regions there is prior knowledge on the distribution of the $K$ components over each one of the $D$ independent observables. Then the data and its mutual information at the event-by-event level provides the information to infer and learn A posterior on the class fractions, and in particular the posterior probability distribution for the signal fraction in the sample. The components distributions over the $D$ observables, whose posteriors are expected to be closer to their true values than its corresponding priors.
Signal should be bounded to a region in phase space, and control regions are needed.
Signal and background can be mixed in different proportions in all phase space. There is no need of a control region.
@@
Figure 2: General Graphical Model for a mixture model. $k$ runs over the $K$ classes, $n$ runs over the $N$ events, and $d$ over the $D$ independent observables. Random variables are represented by circles while arrows represent conditional dependence. White circles represent latent variables which are unobserved while blue circles represent measured variables whose observation conditions the posterior distribution over the parameters. See Ref. [13] for details about representing a probability density function using a Graphical Model.
@@
We assume then exactly two observables ${\cal O}_{1,2}$ and two classes, namely signal ($s$) and background ($b$). In a hard-cut scenario, the observables outputs are simply ${\cal O}_{1}=AB$ or $CD$, and ${\cal O}_{2}=AC$ or $BD$ (see Fig. 1), and thus $p({\bf x}|k)$ is a Categorical distribution for ${\bf x}$ which now consists of four possible results ($A,B,C,D$) and $\theta_{k}$ is a vector of four probabilities which sum up to one. Therefore, we can write
@@ 3 A $hh\to b\bar{b}b\bar{b}$-inspired toy problem
For this toy problem, we assume that the model is correctly specified. That is, that the data follows the model with a specific choice of parameters. Although this will not be a realistic assumption if dealing with physical measurements, this allows us to perform a closure test and more importantly a demonstration of the power of the method. We thus can generate pseudo-data according to a specific choice of parameters of the model and perform Variational Inference (VI) [15] in said pseudo-data to obtain the Maximum a Posteriori (MAP) parameters of the model. A detailed description of the VI algorithm can be found in Ref. [16]. VI is a useful technique that has been implemented in dedicated software such as pyro [17]. We leave a full posterior estimation for future work and consider here just point estimates. This is both numerically easier but also enough if our main goal is to compare directly with the ABCD method. However, one should keep in mind that computing point estimates instead of posterior distributions undersells the power of Bayesian techniques.
@@ 3.1 The toy model for the toy problem
The parameters are sampled from priors with the following hyperparameters
@@
They are chosen to reflect a slight mismodelling of the parameters with reasonable room for the MAP to move. We observe that specific prior hyperparameters are not very important if the number of events is large enough as long as the prior does not forbid the true values. Observe that using priors whose means are shifted from the true values seeks to mimic the expected real case that the Monte Carlo generator is not perfectly tuned to the data,
Figure 4: Data distribution of $b$-tagging score values $\mathcal{S}$ for each of the jet types. True (MAP) distributions are shown in dashed (solid) lines for each jet type, while several distributions sampled from the prior for each individual type are shown in thin solid lines. (Dashed and solid lines have large overlapping.) The MAP distributions are inferred from a dataset with $\pi_{s}=1\%$. The dotted vertical lines correspond to the WP thresholds we use in this work. Notice that data is four-dimensional in the $b$-tagging scores, but here we project it to one-dimension for the sake of showcasing the inference on the individual jet types.
Figure 5: Data distribution of mass values $m$ for each of the individual mass types. True (MAP) distributions are shown in dashed (solid) lines for each mass distribution types, while several distributions sampled from the prior for each type are shown in thin solid lines. The MAP distributions are inferred from a dataset with $\pi_{s}=1\%$. R and NR stand for resonant and non-resonant, respectively.
@@
We generate different datasets with $N=$ 20k total events where $\pi_{s}=0\%$, $0.5\%$, and $1\%$, three reasonable benchmarks where the number of events is high enough for VI to be useful but low enough for finite statistics to provide a relevant limitation. Once inference has been performed, we have access to the MAP for each one of the inferred parameters and therefore to the curves and/or values that they represent in the model. The score and mass generated distributions for a particular dataset can be seen in Figs. 4 and 5, where they are compared to the learned distributions when performing VI.

### v2

@@ 1 Introduction
A more involved, but still brilliantly simple, data-driven modelling technique is the ABCD method [1, 2, 3, 4]. This approach consists in finding two independent observables that classify the background into four regions, such that the signal lies only in one of them. The method allows to predict the signal events in its corresponding region as long as its underlying hypotheses are satisfied. The method has been successfully implemented by the community and it can be found in many of the ATLAS [5] and CMS [6] public results, with some particular examples listed in Refs. [7, 8, 9, 10, 11]. In many applications, the method is adapted or extended to the particular features of the observable and backgrounds. The ABCD approach is quite useful because it reduces the impact of potential biases and inaccuracies in simulations, and because of its simplicity it only relies in finding the observables and regions that fulfill the required assumptions.
@@
In this work we propose to study an existing method from the area of Statistics, more specifically from Bayesian Machine Learning, that consists in disentangling many classes which are mixed in a given dataset. This procedure, known as a mixture model [15], considers that the dataset ${\bf X}$ is a realization of a probability density function for $K$ classes with class probabilities $\pi_{k}$ and per-class probability densities with parameters ${\bf\theta}_{k}$. Using the Bayes Theorem, the method yields –most of the times numerically– a probability distribution for the parameters of the model, $p(\mbox{$\theta$}|{\bf X})$, given a collection of data ${\bf X}$, namely events at LHC, each containing the value of several observables measured in the same event. This distribution is called the posterior of the model. In contrast to ABCD, observables in the mixture model can have continuous distributions that, by definition, contain more information than discrete outputs. A mixture model as presented here assumes that each data point (event) belongs to only one of the possible classes.
The mixture model has among its parameters the expected fraction of each class in the dataset, and in particular for the class that corresponds to the signal, this fraction yields the expected amount of signal in the dataset, which is generally the sought-after unknown. In addition, the Bayesian inference also finds the posterior distribution for all the other parameters as well. This posterior refines the understanding of the physical system, since most of the posterior parameters have a connection to the physics involved in the problem. Moreover, as it can be seen from the mathematical formulation of the mixture model, the number of classes is not fixed (while in the ABCD method there can only be signal and background), and the number of observables of each data point is also not limited to any number (while the ABCD method is restricted to two independent observables). For the purposes of simplicity and performance we also demand independence for the observables measured in each data point in the mixture model.  [^1] A more complex scenario with dependent observables could also be tackled through a mixture model as long as one has some prior-knowledge or clues about this dependence. We do not address this possibility in this work. Another recognizable feature of the mixture models is that they do not need control regions, and that there are no hard cuts to define regions nor hard-assignments to define whether a given event belongs to a specific class. Instead, each event is assigned a probability of belonging to each class.
@@
To compare both methods we use a toy problem whose basic properties are inspired by the di-Higgs production at LHC, with both Higgs states decaying to $b\bar{b}$, namely: $pp\to hh\to b\bar{b}b\bar{b}$. This process, being one of the best candidates to probe the nature of the Higgs boson, and one of the most relevant measurements to be expected by the LHC in the forthcoming years, has been studied in [10, 11]. In these works one can find that the ABCD method with some modifications is utilized as part of the pipeline for extracting the results. The process has several interesting features, such as a small cross-section, large backgrounds and, given the class, the presence of several independent or approximately independent observables, such as the four $b$-tagging scores [^2] There are systematic sources that may yield a small dependence between these scores. Although we ignore this dependence in the toy problem studied along is work, it should be taken into account in a more realistic study. of the jets and the two invariant masses of the corresponding paired jets. As a first step in the analysis of this process with Bayesian inference, we consider in this article a toy model that keeps some of its properties, as well as just two of its backgrounds, and analyse it with the ABCD method and with Bayesian inference. We study and quantify both performances and provide details on why the Bayesian analysis has very good perspectives.
@@ 2.1 The ABCD method: an overview
Where $N_{X}$ refers to the number of events in region $X$. $N_{A}(\mbox{background})$ refers to the number of background events in region $A$. This result is easily understood because if the events are distributed according to the background distribution in which the observables are independent, then one expects the ratio $N_{C}/N_{D}$ to be equal to $N_{A}({\mbox{background}})/N_{B}$. Observe that we assume that only region A contains signal. In the literature one can sometimes find that B is named as control region, and $N_{C}/N_{D}$ as a transfer function.
@@
As it can be appraised, the ABCD method is very clever and simple, and it has worked with excellent results and achievements in HEP, as discussed in the introduction. However, if the hypotheses are not exactly satisfied, the predictions would deviate from their corresponding true values, as expected for any statistical analysis with the same hypotheses. Not only that, but also if the total number of events in B, C or D is small, then its Poisson fluctuation would propagate to the signal and background events predicted in A, a characteristic which is potentially alleviated in our proposed method by using differential information.
Therefore, it is also interesting to analyze some limitations in applying the ABCD method. As a first observation one should notice that the method uses hard cuts to define the regions; whereas mixture models can deal with all the information contained in the dataset. Using this information, applications in statistics have shown improvements when using soft-assignments [16]. That is, each event has a given probability of being signal and a given probability of being background (in mixture models this is usually referred as responsibility). As a second point, one should notice that the ABCD method is limited to two independent observables. Because of this, when there are more independent observables then some of them must be combined to reduce the number of observables, and/or the method is binned in the other observables. In the latter scenario, one computes many ABCD methods according to the binned output of the other observable(s). Also, if it is the case that there are many different backgrounds, then the ABCD method does not have the tools to exploit this knowledge beyond its formula in Eq. 1.
@@ 2.2 Improving and generalizing ABCD through Bayes inference techniques
We model this dataset from the mathematical point of view as the outcome of a probabilistic generative model depicted in the Graphical Model [15] in Fig. 2. Graphical models are a convenient way to depict probabilistic generative models as they allow for clear interpretability but also serve as a bookkeeping technique that can be useful when performing inference. We assume that, given the $n^{th}$ event, its class is determined through the sampling of a categorical latent random variable $\bf{z}_{n}$, which can take $K$ values and is not observed. Then, the $D$ observables in this event are sampled using the distributions from the class corresponding to the value of $\bf{z}_{n}$, and we assume that each observable has no dependence on the other $D-1$ observables once the class has been sampled. Henceforth, although $\bf{z}_{n}$ is not observed, the conditionally independent outcome of the $D$ observables provides information about the probability of the event to belong to each one of the classes. Moreover, the $D$ observables also provide information on their own distributions within each class. Bayesian inference provides a framework to exploit all this data.
In standard inference techniques such as approximate Variational Inference or exact Monte Carlo techniques [15], it is customary to use a 1-of-K representation for the latent variable ${\bf z}_{n}$, expressing the variable as a $K$ dimensional vector indexed with only one non-null entry, when writing down the corresponding probability density associated with the latent variable and the corresponding data point ${\bf x}_{n}$. This implies that the latent variable for a given event is a vector with only one component equal to 1 and all other components 0. With this notation the probability for the $n^{th}$ data point becomes
$p({\bf x}_{n},{\bf z}_{n}|{\bf\theta})=p({\bf z}_{n}|{\bf\pi})\prod_{k=1}^{K}\,\left(\prod_{d=1}^{D}p(x_{nd}|{\bf\theta}_{kd})\,\right)^{z_{n_{k}}},$ (2)
where $z_{n_{k}}$ is the $k^{th}$ component of ${\bf z}_{n}$ and ${\bf\theta}$ is a vector containing all the parameters of the model which can be split into the class fraction parameters ${\bf\pi}=\pi_{k=1,\dots K}$ and the parameters of the distributions for each of the $D$ random variables for the $k$-th class, ${\bf\theta}_{kd}$, which can themselves be a vector depending on the choice of per class and per observable distribution. To shorten the notation, we will refer to the class-dependent distributions simply as $p({\bf x}_{n}|k)$. Having represented ${\bf z}_{n}$ as a 1-of-K vector, we observe how the product selects only the class with a non-zero value of $z_{n_{k}}$. Additionally, $\bf{z}_{n}$ follows a Categorical [^3] A categorical distribution is the usual multinomial distribution specialized to the case where there is only one draw. distribution with the same form
@@
$p({\bf x}_{n}|{\bf\theta})=\sum_{{\bf z}_{n}=\delta_{kk^{\prime}}}p({\bf x}_{n},{\bf z}_{n}|{\bf\theta}),$ (4)
where with $\delta_{kk^{\prime}}$ we mean again a 1-of-K representation, where ${\bf z}_{n}$ is a $K$ dimensional vector indexed by $k^{\prime}$ with only one non-null entry at index $k$. Because each specific value of ${\bf z}_{n}$ selects one class $k$, the marginalized probability distribution takes the usual mixture model form
$p({\bf x}_{n}|{\bf\theta})$ $=$ $\sum_{{\bf z}=\delta_{k,k^{\prime}}}p({\bf x}_{n},{\bf z}_{n}|{\bf\theta})=\sum_{k}p({\bf z}_{n}=\delta_{k,k^{\prime}}|{\bf\pi})\prod_{k^{\prime}=1}^{K}p({\bf x}_{n}|k^{\prime})^{\delta_{k,k^{\prime}}}$ (5)
$=$ $\sum_{k=1}^{K}\prod_{k^{\prime}=1}^{K}\left(\pi_{k^{\prime}}p({\bf x}_{n}|k^{\prime})\right)^{\delta_{k,k^{\prime}}}$
$=$ $\sum_{k=1}^{K}\,\pi_{k}\,p({\bf x}_{n}|k).$
@@
$p({\bf X}|{\bf\theta})=\prod_{n=1}^{N}\,p({\bf x}_{n}|{\bf\theta})$ (6)
which is one of the fundamental components to run Bayesian inference. We need to add a prior probability for the parameters of the model, $p({\bf\theta})$ , and one can in principle obtain the posterior $p({\bf\theta}|{\bf X})$ through Bayes’ theorem
[…]
This usually can be achieved through a variety of numerical techniques, we describe in Section 3 the tools that we have used along this work. The general methodology can then be summarized as a choice of observables ${\bf x}$ and the modelling of the associated likelihood $p({\bf x}|{\bf\theta})$ whose parameters of interest are assigned a prior $p({\bf\theta})$. The method proceeds to extract the posterior distribution, or for simplicity the Maximum a Posteriori (MAP) estimates of the parameters, from the measured unlabeled data ${\bf X}$, which in turn provides us with estimators of the quantities of interest chief among them the expected number of signal events.
@@
2 independent observables. $D$ independent observables, with $D$ arbitrary.
2 components, signal and background. $K$ components can make up the mixture model, the signal and $K-1$ backgrounds. $K$ arbitrary
Prior knowledge on signal and background distributions allows to define the four hard cut regions A, B, C and D, and to assume that the signal is only found in A. Then the method estimates how many signal events are expected in A. Instead of regions there is prior knowledge on the distribution of the $K$ components over each one of the $D$ independent observables. Then the data and its mutual information at the event-by-event level provides the information to infer and learn A posterior on the class fractions, and in particular the posterior probability distribution for the signal fraction in the sample. The components distributions over the $D$ observables, whose posteriors are expected to be closer to their true values than its corresponding priors.
Signal should be bounded to a region in phase space, and control regions are needed. Signal and background can be mixed in different proportions in all phase space. There is no need of a control region.
@@
Figure 2: General Graphical Model for a mixture model. $k$ runs over the $K$ classes, $n$ runs over the $N$ events, and $d$ over the $D$ independent observables. Random variables are represented by circles while arrows represent conditional dependence. White circles represent latent variables which are unobserved while blue circles represent measured variables whose observation conditions the posterior distribution over the parameters. See, for instance, Chapter 8 in Ref. [15] for details about representing a probability density function using a Graphical Model.
@@
We assume then exactly two observables ${\cal O}_{1,2}$ and two classes, namely signal ($s$) and background ($b$). In a hard-cut scenario, the observables outputs are simply ${\cal O}_{1}=AB$ or $CD$, and ${\cal O}_{2}=AC$ or $BD$ (see Fig. 1), and thus $p({\bf x}|k)$ is a Categorical distribution for ${\bf x}$ which now consists of four possible results ($A,B,C,D$) and ${\bf\theta}_{k}$ is a vector of four probabilities which sum up to one. Therefore, we can write
@@ 3 A $hh\to b\bar{b}b\bar{b}$-inspired toy problem
For this toy problem, we assume that the model is correctly specified. That is, that the data follows the model with a specific choice of parameters. Although this will not be a realistic assumption if dealing with physical measurements, this allows us to perform a closure test and more importantly a demonstration of the power of the method. We thus can generate pseudo-data according to a specific choice of parameters of the model and perform Variational Inference (VI) [17] with pyro [18] in said pseudo-data to obtain the Maximum a Posteriori (MAP) parameters of the model. VI learns the MAP parameters by approximating the posterior with a mean field function $q({\bf\theta},{\bf Z})=\delta({\bf\theta}-{\bf\theta}_{\mbox{MAP}})q({\bf Z}|{\bf\theta})$, where ${\bf Z}$ is the set of all hidden latent variables such as class and jet-type assignment. The ${\bf\theta}_{\mbox{MAP}}$ values are obtained by maximizing the Evidence Lower Bound (ELBO) between the posterior proxy $q$ and the joint distribution $p({\bf X},{\bf\theta},{\bf Z})$, which is equivalent to minimizing the Kullback-Leibler divergence between the true posterior and its approximation. When VI is used in this manner, it yields similar results to a global fit with the prior acting as a regulator. However, VI can be used with more general approximation functions that retain the probabilistic nature of the posterior, see e.g. [19]. We leave this full posterior estimation for future work and consider here just point estimates. This is both numerically easier but also enough if our main goal is to compare directly with the ABCD method. However, one should keep in mind that computing point estimates instead of posterior distributions undersells the power of Bayesian techniques.
@@ 3.1 The toy model for the toy problem
As discussed in previous paragraphs, each set of parameters has an associated prior. We have chosen parametric forms for the priors, with associated hyperparameters such that a priori the parameters are distributed as
@@
The priors and their hyperparameters are meant to reflect a slight mismodelling of the parameters with reasonable associated uncertainty. This uncertainty allows us to assess if the algorithm improves our knowledge about the probability of the true values, which is one of our main interests. We have verified that as long as the prior does not forbid the true values, the specific prior hyperparameters are not very impactful because the number of events is sufficiently large. Observe that using priors whose means are shifted from the true values seeks to mimic the expected real case that the Monte Carlo generator is not perfectly tuned to the data,
Figure 4: Top: Data distribution of $b$-tagging score values $\mathcal{S}$ for each of the jet types. True (MAP) distributions are shown in dashed (solid) lines for each jet type, while several distributions sampled from the prior for each individual type are shown in thin solid lines. (Dashed and solid lines have large overlapping.) The MAP distributions are inferred from a dataset with $\pi_{s}=1\%$. The dotted vertical lines correspond to the WP thresholds we use in this work. Notice that data is four-dimensional in the $b$-tagging scores, but here we project it to one-dimension for the sake of showcasing the inference on the individual jet types. Bottom: Difference between MAP and true distributions for each of the jet types.
Figure 5: Top: Data distribution of mass values $m$ for each of the individual mass types. True (MAP) distributions are shown in dashed (solid) lines for each mass distribution types, while several distributions sampled from the prior for each type are shown in thin solid lines. The MAP distributions are inferred from a dataset with $\pi_{s}=1\%$. R and NR stand for resonant and non-resonant, respectively. Bottom: Difference between MAP and true distributions for each of the mass types.
@@
We generate different datasets with $N=$ 20k total events where $\pi_{s}=0\%$, $0.5\%$, and $1\%$, three reasonable benchmarks where the number of events is high enough for VI to be useful but low enough for finite statistics to provide a relevant limitation. Once inference has been performed, we have access to the MAP for each one of the inferred parameters and therefore to the curves and/or values that they represent in the model. The score and mass generated distributions for a particular dataset can be seen in Figs. 4 and 5, where they are compared to the learned distributions when performing VI. We observe how the MAP distribution is an accurate approximation of the true distribution.